# Plasma acceleration of polarized particle beams

### v1

@@ Abstract
Spin-polarized particle beams are of interest for applications like deep-inelastic scattering, e.g. to gain further understanding of the proton’s nuclear structure. With the advent of high-intensity laser facilities, laser-plasma-based accelerators offer a promising alternative to common radiofrequency-based accelerators, as they can shorten the required acceleration length significantly. However, in the scope of spin-polarized particles, they bring unique challenges.
This paper reviews the developments in the field of spin-polarized particles on the basis of the interaction of laser pulses and high-energy particle beams with plasma. The relevant scaling laws for spin-dependent effects in laser-plasma interaction, as well as acceleration schemes for polarized leptons, ions and gamma quanta are discussed.
@@ 1 Introduction
For applications where polarization persistence is a requirement, the dynamics of particle spin have to be thoroughly understood. These applications, e.g., include the areas of high-energy physics and nuclear fusion.
In the field of high-energy physics, polarized particle beams can be utilized for deep-inelastic scattering to improve the understanding of the proton nuclear structure [6] and are of general relevance to quantum chromodynamics [7]. For nuclear fusion, Goldhaber et al. proposed in [8] that the relevant cross-section depends on the relative spin orientation of the reactants. For the reaction $d+t\to\alpha+n$, the enhancement factor is approximately 50% [9]. For other reactions, like proton-Boron fusion, similar factors have been calculated [10]. Moreover, the understanding of spin effects in plasma has led to the idea of using polarization as a diagnostic tool for measuring the strength of electromagnetic fields in astrophysical plasma [11, 12].
@@
This review provides an overview of the progress in the acceleration of spin-polarized particle beams via plasma-based acceleration schemes. The remainder of the introduction discusses the main spin-dependent effects that can become relevant during of high-intensity laser pulses or high-energy particle beams with plasma. Section 2 describes some of the experimental techniques involved with spin, ranging from the production of polarized particle beams to their detection. Results on polarized electrons and positrons are presented in section 3. Similarly, theoretical and experimental results on polarized ions are presented in 4. The production of polarized gamma quanta is described in 5. Lastly, a summary and a discussion of future prospects are found in section 6.
@@
An extensive discussion of the effects relevant to laser-plasma based sources of polarized beams was given by Thomas et al. in [13], in which the main effects discussed are spin precession, spin-dependent trajectories and radiative polarization. A schematic overview of the various effects’ connections is presented in Figure 1. In the following, we will detail the most relevant aspects of this discussion.
@@
Figure 1: Schematic overview of the relevant effects affecting particle spin in laser-plasma interaction. Reproduced under the terms of the CC-BY license from [13]. Copyright 2020, The Authors, published by American Physical Society.
@@
[…]
@@
    depend on the anomalous magnetic moment $a$ and the Lorentz factor $\gamma$ of the corresponding particle. The anomalous magnetic moment is defined as $a=(g-2)/2$ via the Landé $g$-factor. It can be calculated as higher-order loop corrections to the magnetic moment. For the electron it is $a_{e}\approx\alpha/(2\pi)\approx 10^{-3}$, for protons it is roughly $m_{p}/m_{e}$ times larger.
@@
    where $P_{\uparrow},P_{\downarrow}$ denote the different spin-flip probabilities.
@@
More extensive discussions on radiative spin dynamics have been given by several authors in the framework of QED, like studies on spin-dependent radiation reaction by Seipt et al. [18]. Geng et al. identified a spin-dependent deflection mechanism based on radiation reaction that occurs during laser-electron interaction [19]. The effect is shown to dominate the Stern-Gerlach force for electrons with 100s of MeV and 10 PW-class laser pulses. Over the years, various models describing radiative polarization have been proposed [20, 21, 22, 23, 24], which are discussed in more detail in the sections concerning polarized positrons and gamma quanta. A kinetic description for relativistic spin-polarized plasma was derived by Seipt and Thomas in [25]. Therein, QED effects are taken into account via Boltzmann-type collision operators and the approximation of locally constant fields.
@@ 2.1.1 Pre-polarized targets
The only polarized target that has been used so far for laser-plasma acceleration experimentally employs a jet containing nuclear polarized 3He gas [29]. This rare helium isotope has the advantage that it can be rather easily polarized at room temperature and be stored for many hours in a magnetic holding field (cf. Fig. 2). High degrees of nuclear polarization ($P\sim 40$%) can be obtained in a gas jet at densities of a few times $10^{19}$ cm-3. The polarization for this and similar targets can be achieved via methods like metastability-exchange optical pumping [30] or spin-exchange optical pumping [31]. Data of an experiment at the PHELIX laser suggest persistence of this polarization after acceleration of the 3He ions to MeV energies [32]. A disadvantage of pre-polarized targets, such as 3He gas, is that a permanent magnetic holding field must be provided around the interaction zone in order to maintain the target polarization. In the case of 3He gas, the field strength requirement is rather low, on the order of a few mT, but great care has to be taken on the homogeneity of the holding field [29]. Such fields can either be provided by permanent-magnet assemblies or by a Helmholtz-type coils. These have the disadvantage of requiring cooling (if mounted inside the laser-interaction chamber) but allow the initial ${}^{3}\mkern-2.0mu$e polarization to be rotated relative to the laser propagation axis, thus providing maximum flexibility in adjusting the beam polarization.
@@
While the use of SPH targets seems rather straightforward for proton acceleration [35, 36], the acceleration of polarized electrons would first require the removal of unpolarized electrons from the accelerating plasma which are unavoidably present in the (otherwise unused) halide atoms. Two solutions have been put forward here, either removal via ionization [37] or spatial separation of the SPH from the unwanted photoproducts leaving behind only hydrogen atoms of high polarization ($P=90$%) [34]. It remains to be seen which of these approaches will ultimately prove to be more practicable. A schematic detailing the production of polarized electrons from an HCl target is presented in Fig. 3.
According to Ref. [38], hyperpolarized cryogenic foils can be produced after the recombination of nuclear polarized hydrogen atoms from an atomic beam source to hydrogen-deuterium molecules. These could in principle be used as targets for plasma acceleration of polarized proton or deuterium beams. Due to their technically demanding realization [39], it has yet to be shown that this type of target is suitable for laser-plasma applications. A theoretical study of the polarization dynamics in the hydrogen isotope molecules, H2, D2, and HD can be found in Ref. [40]. It could be shown that in the presence of a static magnetic field the molecular rotation polarization can be transferred and maintained in the nuclear spin, providing highly nuclear-spin-polarized molecules ($P>90$%).
@@ 2.1.2 Unpolarized targets
An unpolarized foil target was used for the first polarization measurement of laser-accelerated particles. The authors of Ref. [42] report on the TNSA acceleration of few-MeV protons from the rear surface of a gold foil. Since no polarization build-up of the protons during the acceleration was observed, it was concluded that pre-polarized targets would be needed to produced polarized hadron beams.
@@
Figure 2: Magnetic holding setup for the acceleration of polarized Helium-3. Reproduced under the terms of the CC-BY license from [29]. Copyright 2022, The Authors, published by MDPI AG.
Figure 3: (a) Potential experimental configuration for LWFA of spin-polarized electrons. (b) Schematic of polarized electron production from an HCl target using multiple lasers for alignment, photo dissociation and ionization. Reproduced under the terms of the CC-BY license from [37]. Copyright 2019, The Authors, published by IOP Publishing Ltd.
@@ 2.2 Polarimetry
A compact polarimeter for 3He ions with special emphasis on the analysis of short-pulsed beams accelerated during laser–plasma interactions is described in Ref. [45]. The polarimeter was tested at a tandem accelerator with unpolarized 3He beams and then used for the analysis of polarized laser-accelerated ions [32]. It is based on the detection of $\alpha$ particles in CR-39 stacks, emitted from the $d$-3He fusion reaction with non-vanishing $A_{y}$ in a deuterated-polyethylene foil. The polarimeter is especially useful for the analysis of 3He ion beams in the few-MeV range, but also at lower energies down to 0.6 MeV. The accuracy of a beam polarization measurement, i.e., the minimal detectable 3He beam polarization, is estimated to be 21% for a typical laser–plasma experiment where roughly $10^{9}$ 3He ions enter the polarimeter.
@@ 3 Polarized electrons and positrons
Within this section, we transition to the results obtained in – up to now – only theoretical studies of the acceleration of polarized lepton beams. Most of the results for electrons are concerned with wakefield acceleration. As mentioned in the previous section, most of the following setups are based on pre-polarized targets, and only a few make use of “single-step” schemes to polarize and directly accelerate the particles.
@@ 3.1 Electrons
Further studies on spin precession in wakefields were performed by Pugacheva et al. in [48]. It was shown that the minimal depolarization occurs for electrons that are injected close to the maximum of the accelerating force with the wake’s phase velocity. In a follow-up publication [49], the effects of synchrotron radiation on spin dynamics was discussed.
@@
An alternative means of injection has been discussed in the form of colliding-pulse injection. Gong et al. proposed that the driving laser pulse collides with a weaker laser pulse, which leads to a heat-up of electrons in the region of their interaction [54]. Due to this heating, electrons can be injected into the wake. The mechanism is examined for the polarized HCl target at $10^{18}$ cm-3. In [55], the driving pulse is fixed at $a_{0}=2.5$ and $w_{0}=20$ $\mu$m. The intensity and focal spot size of the laser pulse are varied, examining the influence on beam charge and polarization. It is found that similar spot size between driving and colliding pulse, as well as higher pulse intensity lead to an increase in charge, however at the cost of polarization (cmp. Fig. 4). Electron beams with 80% polarization with tens of pC can be generated using 10 TW-class laser systems; beams in excess of 90% polarization are achievable for lower charge.
Figure 4: Dependence of beam charge and polarization degree, as well as emittance and peak current, on parameters of the colliding laser pulse. The driving laser pulse is fixed to $a_{0}=2.5$ and $w_{0}=20$ $\mu$m. Reproduced under the terms of the CC-BY license from [55]. Copyright 2023, The Authors, published by American Physical Society.
@@
The case of beam-driven wakefields was first discussed by Wu et al. in [57]. Therein, the influence of driving beam parameters on the witness are studied. Analytical scaling laws for beam polarization in dependence of the peak current are derived, which show that polarization decreases for increases flux. Up to 7.5 kA of electrons with 80% polarization are obtained from simulations, which is comparable to their LWFA results with respect to the polarization degree.
@@
The aforementioned studies for laser- and beam-driven wakefields generally assume a pre-polarized target, meaning that target preparation and acceleration is a two-step process. Nie et al., however, proposed methods to achieve polarized electron beams in-situ [43]. The electrons are injected into the wakefield and simultaneously polarized via the ionization of the outermost $p$ orbital of the utilized noble gas. In the publication, a target consisting of lithium and xenon is used. An electron drive beam ionizes only the lithium and generates the wakefield. A circularly polarized laser pulse is then used to strong-field ionize the $5p^{6}$ electron of the xenon atoms. These electrons are spin-polarized and can be accelerated to approx. 2.7 GeV in 11 cm. A net polarization of about 31% is achieved with this setup. In a subsequent publication, Nie et al. investigate the potential of single-species (ytterbium) plasma photocathodes for polarized beams [44]. Here, the electron beam predominantly ionizes the outer $6s$ electrons and drives the wakefield. A CP laser pulse is then used to inject the $4f^{14}$ electrons which is shown to deliver 56% net polarization with 15 GeV in 41 cm. The setup is shown in Figure 5.
@@
Figure 5: The single-species photocathode scheme for polarized electrons. (a), (b) show the plasma and beam density around the laser focus and after driver pinching, respectively. (c), (d) show the Yb ion density at the corresponding time steps. Reproduced under the terms of the CC-BY license from [44]. Copyright 2022, The Authors, published by American Physical Society.
@@
Beyond the concept of wakefield acceleration from pre- or unpolarized targets in the setting of classical physics, other studies have focused on the radiative polarization dynamics of electrons in intense electromagnetic fields, like Tang et al. [23]. Seipt et al. proposed the polarization of high-energy electron beams using pulsed bi-chromatic laser fields [18]. As a potential application, Gong et al. showed that initially unpolarized targets of near-critical density can be used as a diagnostic tool for transient magnetic fields [11]. Similarly, the radiative polarization of electrons during current filamentation instability offers a novel pathway to study astrophysical plasma [12]. Further references on radiative polarization can be found in the introductory part of this review as well as the section concerning polarized gamma quanta.

### v2

@@ Abstract
Spin-polarized particle beams are of interest for applications like deep-inelastic scattering, e.g. to gain further understanding of the proton’s nuclear structure. With the advent of high-intensity laser facilities, laser-plasma-based accelerators offer a promising alternative to standard radiofrequency-based accelerators, as they can shorten the required acceleration length significantly. However, in the scope of spin-polarized particles, they bring unique challenges.
This paper reviews the developments in the field of spin-polarized particles, focusing on the interaction of laser pulses and high-energy particle beams with plasma. The relevant scaling laws for spin-dependent effects in laser-plasma interaction, as well as acceleration schemes for polarized leptons, ions, and gamma quanta, are discussed.
@@ 1 Introduction
For applications in areas such as high-energy physics and nuclear fusion, where polarization persistence is a requirement, the dynamics of particle spin have to be thoroughly understood.
In the field of high-energy physics, polarized particle beams can be utilized for deep-inelastic scattering (DIS) to improve the understanding of the proton nuclear structure [6] and are of general relevance to quantum chromodynamics [7]. Unpolarized DIS gives insights into the number density of partons with a momentum fraction $x$ of the hadron. By contrast, polarized DIS, i.e. the collision of a longitudinally polarized lepton beam with a longitudinally or transversely polarized target, yields information on the number density of partons for specific $x$ and spin polarization [8]. The spin-dependent scaling functions $g_{1,2}(x)$ are found to have more complicated interpretations in a naive parton model of the nucleons than the unpolarized equivalents $F_{1,2}(x)$. Thus, the polarized process has been used to determine to what percentage the proton spin is carried by the quarks [9]. Similarly, polarized Drell-Yan processes can also be used to gain insight into parton distribution functions [10].
For nuclear fusion, Goldhaber et al. proposed in [11] that the relevant cross section depends on the relative spin orientation of the reactants. For the reaction $d+t\to\alpha+n$, the enhancement factor is approximately 50% [12]. The intermediate step in the reaction, the nucleus $^{5}\mathrm{He}$, has an excited state with angular momentum $J=3/2$. The spins of the deuteron and triton are $1\hbar$ and $\frac{1}{2}\hbar$, respectively, thus leading to a total angular momentum of $J\hbar$, with $J\in\{\frac{1}{2},\frac{3}{2}\}$. Since the statistical weight for the $3/2$-state is four, and it is two for the $1/2$-state, random polarization of both reactants yields a probability of $2/3$ that the colliding system ends up in the state of $J=3/2$ [13]. Thus, if the spins are aligned with respect to the magnetic field, the cross section is increased, yielding a high probability that the excited $^{5}\mathrm{He}$ nucleus is formed, with subsequent decay into $\alpha+n$. For other reactions, like proton-boron fusion, similar factors have been calculated [14].
Low-energy, polarized electrons in the 100 eV range are also of interest for applications in surface physics, including research on spin-orbit coupling and topological materials [15]. This has brought forward techniques like spin-polarized low-energy electron diffraction (SPLEED) to study scattering from single-crystal solid surfaces [16].
Moreover, the understanding of spin effects in plasma has led to the idea of using polarization as a diagnostic tool for measuring the strength of electromagnetic fields in astrophysical plasma [17, 18] or for examining the structure of plasma wakefields [19].
The prospects of these applications have led to increased research on sources of polarized particle beams and their plasma-based acceleration. Beyond that, polarimetry tailored to these novel sources has been developed in recent years, as the time structure of the particle beams and the strong electromagnetic pulses induced by the accelerating laser are challenging for conventional detectors.
@@
This review provides an overview of the progress in the acceleration of spin-polarized particle beams via plasma-based acceleration schemes. The remainder of the introduction discusses the main spin-dependent effects that can become relevant during the interaction of high-intensity laser pulses or high-energy particle beams with plasma. Section 2 describes some of the experimental techniques involved with spin, ranging from the production of polarized particle beams to their detection. The results on polarized electrons and positrons are presented in section 3. Similarly, theoretical and experimental results on polarized ions are presented in 4. The production of polarized gamma quanta is described in 5. Lastly, a summary and a discussion of future prospects are found in section 6.
@@
An extensive discussion of the effects relevant to laser-plasma based sources of polarized beams was given by Thomas et al. in [20], in which the main effects discussed are spin precession, spin-dependent trajectories, and radiative polarization. A schematic overview of the connection of the various effects is presented in Figure 1. In the following, we will detail the most relevant aspects of this discussion.
@@
Figure 1: Schematic overview of the relevant effects affecting particle spin in laser-plasma interaction. Reproduced under the terms of the $C$-BY license from [20]. Copyright 2020, The Authors, published by American Physical Society.
@@
    is the precession frequency. It depends on the prevalent electromagnetic fields $\textbf{E},\textbf{B}$ and the particle’s velocity v. Further, $qe$ denotes the charge of the particle, $m$ its mass and $c$ is the vacuum speed of light. The prefactors
@@
    depend on the anomalous magnetic moment $a$ and the Lorentz factor $\gamma$ of the corresponding particle. The anomalous magnetic moment is defined as a=(g2)/2a=(g-2)/2 via the Landé $g$-factor. It can be calculated as higher-order loop corrections to the magnetic moment. For the electron it is $a_{e}\approx\alpha/(2\pi)\approx 10^{-3}$, for protons it is roughly $m_{p}/m_{e}$ times larger. An extensive list of different magnetic dipole moments is given, e.g., by Stone [23].
@@
    where $P_{\uparrow},P_{\downarrow}$ denote the different spin-flip probabilities and $\tau_{\mathrm{pol}}$ is the polarization time.
@@
More extensive discussions on radiative spin dynamics have been given by several authors in the framework of quantum electrodynamics (QED), like studies on spin-dependent radiation reaction by Seipt et al. [26]. Geng et al. identified a spin-dependent deflection mechanism based on radiation reaction that occurs during the laser-electron interaction [27]. The effect is shown to dominate the Stern-Gerlach force for electrons with 100s of MeV and 10 PW-class laser pulses. Over the years, various models describing radiative polarization have been proposed [28, 29, 30, 31, 32, 33], which are discussed in more detail in the sections concerning polarized positrons and gamma quanta. A kinetic description of relativistic spin-polarized plasma was derived by Seipt and Thomas in [34]. Therein, QED effects are taken into account via Boltzmann-type collision operators and the approximation of locally constant fields.
@@ 2.1.1 Pre-polarized targets
The only polarized target that has been used so far for laser-plasma acceleration experimentally employs a jet containing nuclear polarized 3He gas [38]. This rare helium isotope has the advantage that it can be rather easily polarized at room temperature and be stored for many hours in a magnetic holding field (cf. Fig. 2). High degrees of nuclear polarization ($P\sim 40$%) can be obtained in a gas jet at densities of a few times $10^{19}$ cm-3. The polarization for this and similar targets can be achieved via methods like metastability-exchange optical pumping [39] or spin-exchange optical pumping [40]. Data of an experiment at the PHELIX laser suggest persistence of this polarization after acceleration of the 3He ions to MeV energies [41]. A disadvantage of pre-polarized targets, such as 3He gas, is that a permanent magnetic holding field must be provided around the interaction zone in order to maintain the target polarization. In the case of 3He gas, the field strength requirement is rather low, on the order of a few mT, but great care has to be taken with the homogeneity of the holding field [38]. Such fields can either be provided by permanent-magnet assemblies or by Helmholtz-type coils. The latter have the disadvantage of requiring cooling (if mounted inside the laser-interaction chamber), but allow the initial $^{3}\mkern-2.0mu$He polarization to be rotated relative to the laser propagation axis, thus providing maximum flexibility in adjusting the beam polarization.
@@
While the use of SPH targets seems rather straightforward for proton acceleration [44, 45], the acceleration of polarized electrons would first require the removal of unpolarized electrons from the accelerating plasma which are unavoidably present in the (otherwise unused) halide atoms. Two solutions have been put forward here, either removal via ionization [46] or spatial separation of the SPH from the unwanted photoproducts leaving behind only hydrogen atoms of high polarization (P=90P=90%) [43]. It remains to be seen which of these approaches will ultimately prove to be more practicable. A schematic detailing the production of polarized electrons from an HCl target is presented in Fig. 3.
Kannis and Rakitzis have proposed an alternative using formaldehyde or formic acid [47]. These molecules can be excited using infrared laser light. The gained rotational polarization can be subsequently transferred to nuclear polarization via hyperfine polarization beating. Afterwards, the molecules are photodissociated to the target molecules $\mathrm{H}_{2}$ or $\mathrm{H}_{2}\mathrm{O}$. The production rates of this mechanism can approach $10^{21}\;\mathrm{s}^{-1}$, in contrast to molecular-beam separation techniques which deliver rates of only up to $10^{12}\;\mathrm{s}^{-1}$. This is of importance of potential applications in nuclear magnetic resonance and polarized fusion tests, which require rates of at least $10^{17}\;\mathrm{s}^{-1}$.
According to Ref. [48], hyperpolarized cryogenic foils can be produced after recombination of nuclear polarized hydrogen atoms from an atomic beam source to hydrogen-deuterium molecules. These could in principle be used as targets for plasma acceleration of polarized proton or deuterium beams. Due to their technically demanding realization [49], it has yet to be shown that this type of target is suitable for laser-plasma applications. A theoretical study of the polarization dynamics in the hydrogen isotope molecules, H2, D2, and HD can be found in Ref. [50]. It could be shown that in the presence of a static magnetic field the molecular rotation polarization can be transferred and maintained in the nuclear spin, providing highly nuclear-spin-polarized molecules (P>90P>90%).
@@ 2.1.2 Unpolarized targets
An unpolarized foil target was used for the first polarization measurement of laser-accelerated particles. The authors of Ref. [52] report on the TNSA acceleration of few-MeV protons from the rear surface of a gold foil. Since no polarization build-up of the protons during the acceleration was observed, it was concluded that pre-polarized targets would be needed to produce polarized hadron beams.
@@
Figure 2: Magnetic holding setup for the acceleration of polarized Helium-3. Reproduced under the terms of the $C$-BY license from [38]. Copyright 2022, The Authors, published by MDPI AG.
Figure 3: (a) Potential experimental configuration for laser-driven wakefield acceleration of spin-polarized electrons. (b) Schematic of polarized electron production from an HCl target using multiple lasers for alignment, photo dissociation and ionization. Reproduced under the terms of the $C$-BY license from [46]. Copyright 2019, The Authors, published by IOP Publishing Ltd.
@@ 2.2 Polarimetry
A compact polarimeter for 3He ions with special emphasis on the analysis of short-pulsed beams accelerated during laser–plasma interactions is described in Ref. [55]. The polarimeter was tested at a tandem accelerator with unpolarized 3He beams and then used for the analysis of polarized laser-accelerated ions [41]. It is based on the detection of $\alpha$ particles in CR-39 stacks, emitted from the dd-3He fusion reaction with non-vanishing $A_{y}$ in a deuterated-polyethylene foil. The polarimeter is especially useful for the analysis of 3He ion beams in the few-MeV range, but also at lower energies down to 0.6 MeV. The accuracy of a beam polarization measurement, i.e., the minimal detectable 3He beam polarization, is estimated to be 21% for a typical laser–plasma experiment where roughly $10^{9}$ 3He ions enter the polarimeter.
@@
Laser-Compton polarimetry usually refers to Compton back-scattering polarimeters [57], measuring a longitudinally polarized electron beam at high energies ($\gtrsim 1$ GeV) via the collision with circularly polarized laser light (532 nm, photon energy 2.33 eV), which produces backscattered photons ($\gtrsim 10$ MeV) [58] with a rate depending on the the spin-dependent Compton cross section. The applicability of this method to electron beams from laser-driven wakefield acceleration (LWFA) is very limited, since it would require a precisely calibrated beam-intensity monitor to be operated together with the Compton polarimeter. A non-destructive diagnostic method for LWFA electron bunches based on Transition Radiation (TR) is being studied [59], but the technique is still under development [60, 61] with a comparison to the standard destructive method for energy and charge measurement by a magnetic spectrometer.
Compton transmission polarimetry has been performed for a polarized electron beam of 3.5 MeV at the MAMI injector [62], a longitudinally polarized electron beam ($P_{e}\sim 80\%$) at the A4 MAMI experiment at 570 MeV and 854 MeV [63], and also as relative electron polarization monitors at 20 and 200 MeV at MIT Bates [64], or for polarized positrons (selected energies from 4.5 to 7.5 MeV) and electrons (6.7 MeV) with undulator-based production from E166 at SLAC [65]. A recent application is for DC and SRF electron photo-injectors at JLab and BNL, i.e., spin-polarized electron beam sources at 5 and 7  MeV [66]. Compton transmission polarimeters have several advantages for LWFA electron bunches. First, they can be applied for electron (or positron and $\gamma$-photon) bunches with any time structure [67], which is crucial, since the duration of the bunches is generally less than 10 fs and the number of particles is about $10^{9}$ ($\sim$0.1 nC, and the equivalent current $\sim$10 kA), ruling out alternative methods. Second, the Compton transmission polarimeters can be used for non-destructive measurements at high energies since the electrons only loose a small fraction of their energy (below a few MeV, also keeping most of their polarization) after a thin radiator and can be separated by a sweeping magnet [64]. This method allows measuring the beam polarization and the beam energy (with an additional electron spectrometer) simultaneously. The beam profile and energy distribution have shot-to-shot fluctuations (laser-plasma instabilities [2, 68]) in LWFA, so the correction must be included in the polarization measurement. Finally, it is suitable for very high beam intensities because the transmitted $\gamma$-rays are collimated to a small forward cone and reduced by the magnetized absorber and end in a calorimeter with a large energy capacity. As discussed, several of the aforementioned polarimetry techniques have been successfully demonstrated. However, adapting them to the developing array of sources for leptons, ions, and gamma quanta described in the subsequent chapters remains an on-going challenge.
@@ 3 Polarized electrons and positrons
Within this section, we transition to the theoretical results obtained in studies of the acceleration of polarized lepton beams. Most of the results for electrons are concerned with wakefield acceleration. As mentioned in the previous section, most of the following setups are based on pre-polarized targets, and only a few make use of “single-step” schemes to polarize and directly accelerate the particles.
@@ 3.1 Electrons
Further studies on spin precession in wakefields were performed by Pugacheva et al. in [70]. It was shown that the minimal depolarization occurs for electrons that are injected close to the maximum of the accelerating force with the wake’s phase velocity. In a follow-up publication [71], the effects of synchrotron radiation on spin dynamics were discussed.
@@
[…]
@@
The case of beam-driven wakefields was first discussed by Wu et al. in [79]. Therein, the influence of driving beam parameters on the witness are studied. Analytical scaling laws for beam polarization in dependence of the peak current are derived, which show that polarization decreases for increases flux. Up to 7.5 kA of electrons with 80% polarization are obtained from simulations, which is comparable to their LWFA results with respect to the polarization degree. Beam-driven wakefields can have several advantages over LWFA as effects like laser diffraction and dephasing do not occur. However, beam-driven approaches require a particle source which is not readily available when compared to laser sources, and other detrimental effects like the hosing instability have to be mitigated depending on the specific parameter regime.
For the approaches using pre-polarized targets in connection wakefield acceleration, it should be noted that many publications consider a background density in the range of $10^{18}\;\mathrm{cm}^{-3}$, whereas experimentally feasible densities are currently in the regime of $\leq 10^{17}\;\mathrm{cm}^{-3}$ [43]. Moreover, the volume containing polarized particles that is available for laser-plasma interaction is limited, restricting the injection/acceleration methods for polarized electrons significantly. This is due to method with which the hydrogen halide targets are prepared: the second dissociation laser pulse is responsible for shaping the volume in which only the polarized hydrogen is located. Since collisions which would otherwise decrease the polarization degree have to be avoided, the maximum attainable density is limited. In the case of [43], the HCl target is modeled as a channel of 20 µm width and a density of $10^{17}\;\mathrm{cm}^{-3}$. The laser pulse driving the acceleration process has a wavelength of 1.6 µm and $a_{0}=6$. An electron beam with 3.9 pC can be produced that way, maintaining 48% of its initial polarization. As the available degree of pre-polarization depends on channel width and target density (here: 45 %), this means that effectively only 22% polarization can be achieved based on these target restrictions.
@@
The aforementioned studies for laser- and beam-driven wakefields generally assume a pre-polarized target, meaning that target preparation and acceleration is a two-step process. Nie et al., however, proposed methods to achieve polarized electron beams in-situ [53]. The electrons are injected into the wakefield and simultaneously polarized via the ionization of the outermost $p$-orbital of the utilized noble gas. In the publication, a target consisting of lithium and xenon is used. An electron drive beam ionizes only the lithium and generates the wakefield. A circularly polarized laser pulse is then used to strong-field ionize the $5p^{6}$ electron of the xenon atoms. These electrons are spin-polarized and can be accelerated to approx. 2.7 GeV in 11 cm. A net polarization of about 31% is achieved with this setup. In a subsequent publication, Nie et al. investigate the potential of single-species (ytterbium) plasma photocathodes for polarized beams [54]. Here, the electron beam predominantly ionizes the outer $6s$ electrons and drives the wakefield. A circularly polarized (CP) laser pulse is then used to inject the $4f^{14}$ electrons which is shown to deliver 56% net polarization with 15 GeV in 41 cm. The setup is shown in Figure 5.
@@
Figure 5: The single-species photocathode scheme for polarized electrons. (a), (b) show the plasma and beam density around the laser focus and after driver pinching, respectively. (c), (d) show the Yb ion density at the corresponding time steps. Reproduced under the terms of the $C$-BY license from [54]. Copyright 2022, The Authors, published by American Physical Society.
Beyond the concept of wakefield acceleration from pre- or unpolarized targets in the setting of classical physics, other studies have focused on the radiative polarization dynamics of electrons in intense electromagnetic fields, like Tang et al. who propose a self-consistent model that extends the applicability of the polarization vector model to electromagnetic field structures relevant to laser-plasma interaction [31]. Zhu et al. follow an approach utilizing beam-solid interaction [80]: here, a relativistic, unpolarized electron beam is incident onto a solid target at a grazing angle. This leads to strong magnetic fields and, in turn, strong beam self-focusing via magnetic pinching. The magnetic fields reach the level of a few Giga-Gauss, triggering synchrotron radiation and accordingly, radiative spin flips. In their simulations, they find that an initial 5 GeV electron beam can generate 50% polarization for electrons with <2<2 GeV using their approach. On a similar note, Xue et al. utilize a double-layer solid target which is ionized by the incident electron beam [81]. The ionization leads to an asymmetric plasma field which can induce spontaneous radiative polarization. This method has the added benefit that the generated fields can further focus the polarized beam. In their 3D-PIC simulations, they are able to produce pC-level electron beams with a polarization of approx. 57%. While these approaches provide high polarization and energy of the resulting electron beams, they also require near-future laser facilities for realization. In contrast, Seipt et al. proposed the polarization of high-energy electron beams using pulsed bi-chromatic laser fields, which could be realized with current-day parameters, but this delivers a lower beam polarization of approx. 17% [26].
@@
As a potential application, Gong et al. showed that initially unpolarized targets of near-critical density can be used as a diagnostic tool for transient magnetic fields [17]. They find that the angle-resolved electron spin polarization can be used to be predict the leading order of the quasi-static magnetic field induced by the longitudinal current. The method is shown to be robust to moderate imperfections of both laser pulse and plasma density profile. Similarly, An et al. proposed to utilize polarized electron beams to reconstruct the structure of a wakefield, providing advantages compared to conventional charged particle beam probes for more complex structures [19]. Further, the radiative polarization of electrons during current filamentation instability has been shown to offer a novel pathway to study astrophysical plasma, allowing to differentiate between different instability regimes like (ab-)normal filamentation and quenching [18]. Further references on radiative polarization can be found in the introductory part of this review as well as the section concerning polarized gamma quanta.